# AnomalyGFM: Graph Foundation Model for Zero/Few-shot Anomaly Detection

### v1

@@ Appendix A Detailed Dataset Description
  • Amazon (Dou et al., 2020): It includes product reviews under the Musical Instrument category. The users with more than 80% of helpful votes were labeled as begin entities, with the users with less than 20% of helpful votes treated as fraudulent entities. There are three relations including U-P-U (users reviewing at least one same product), U-S-U (users giving at least one same star rating within one week), and U-V-U (users with top-5% mutual review similarities). In this article, we do not distinguish this connection and regard them as the same type of edges, i.e., all connections are used. There are 25 handcrafted features that were collected as the raw node features.
@@
  • Amazon-all (Dou et al., 2020): It includes three types of relations: U-P-U (users reviewing at least one same product), U-S-U (users giving at least one same star rating within one week), and U-V-U (users with top-5% mutual review similarities). Amazon-all is formed by treating the different relations as a single relation following (Chen et al., 2022; Qiao and Pang, 2023a).
  • YelpChi-all (Dou et al., 2020): Similar to Amazon-all, YelpChi-all includes three types of edges: R-U-R (reviews posted by the same user), R-S-R (reviews for the same product with the same star rating), and R-T-R (reviews for the same product posted in the same month). YelpChi-all is formed by treating the different relations as a single relation following (Chen et al., 2022; Qiao and Pang, 2023a).
@@
  • Question (Platonov et al., 2023): It is collected from the website Yandex Q where the node represents the user and the edge connecting the node represents the user who has answered other’s questions during a one year period. The attribute of nodes is the mean of FastText embeddings for words in the description of the user. For the user without description, the additional binary features are employed as the feature of the user.
  • Elliptic (Weber et al., 2019): It is a Bitcoin transaction network in which the node represents the transactions and the edge is the flow of Bitcoin currency. The Bitcoin transaction is mapped to real-world entities associated with licit categories in this dataset.
  • T-Finance (Tang et al., 2022): It is a financial transaction network where the node represents an anonymous account and the edge represents two accounts that have transaction records. The features of each account are related to some attributes of logging like registration days, logging activities, and interaction frequency, etc. The users are labeled as anomalies when they fall into the categories like fraud money laundering and online gambling.
@@
Appendix B Description of Baselines
B.1. Competing GAD Methods
A more detailed introduction of the eight GAD models we compare with is given as follows.
Table 6. The comparison of AUROC and AUPRC results on 8 real-world GAD datasets under zero-shot setting with the models trained on Amazon only. For each dataset, the best performance per column within each metric is boldfaced, with the second-best underlined. All results are the average of three runs with different random seeds.
@@ B.1. Competing GAD Methods
  • AnomalyDAE (Fan et al., 2020) consists of a structure autoencoder and an attribute autoencoder to learn both node embeddings and attribute embeddings jointly in a latent space. In addition, an attention mechanism is employed in the structure encoder to capture normal structural patterns more effectively.
  • CoLA (Liu et al., 2021b) exploits the local information of the network and introduces the contrastive self-supervised learning framework for GAD which can capture the relationship between the target node and its neighbors. The abnormality is reflected by measuring the similarity of the node and its contextual nodes.
  • GADAM (Chen et al., 2024) introduces a hybrid attention-based adaptive message-passing mechanism, leveraging local inconsistency mining to selectively incorporate both normal and abnormal signals from contextual nodes.
  • TAM (Qiao and Pang, 2023b) learns tailored node representations for a new anomaly measure by maximizing the local affinity of nodes to their neighbors. TAM is optimized on truncated graphs where non-homophily edges are removed iteratively. The learned representations result in a significantly stronger local affinity for normal nodes than abnormal nodes. So, a local affinity-based anomaly measure is used to calculate the anomaly score.
  • GCN (Kipf and Welling, 2016) performs neighborhood aggregation (or message passing), where each node updates its representation by aggregating the features of its neighbors.
  • GAT (Veličković et al., 2017) introduces attention mechanisms to GCN by assigning learnable attention weights to each edge, allowing the model to focus more on important neighbors.
  • BWGNN (Tang et al., 2022) reveals the ’right shift’ phenomenon that the spectral energy distribution concentrates more on the node with high frequencies and less on the node with low frequencies and proposes a spectral and spatial localized band-pass filter to address this problem to achieve better anomaly detection.
  • GHRN (Gao et al., 2023) first indicates that heterophily is positively associated with spectral frequencies in the graph. To fully exploit this observation, they prune the inter-class edge to emphasize and delineate the high-frequency components of the graph, thereby improving the performance of the binary classifier.
B.2. Competing Generalist Methods
  • GraphPrompt (Liu et al., 2023) builds a pre-training and prompting framework designed for both graph-level and node-level classification tasks. It learns a prompt that can help the downstream tasks to identify and leverage the most relevant knowledge from the pre-trained model in a task-specific way.
  • ARC (Liu et al., 2024) is the first fine-tuning generalist GAD approach that enables a “one-for-all” GAD model. It proposes an ego-neighbor residual graph encoder that learns abnormality-related node embedding. In the inference step, a cross-attentive in-context anomaly scoring module is employed to predict node abnormality by leveraging few-shot normal samples.
  • UNPrompt (Niu et al., 2024) is the first zero-shot generalist GAD method that learns generalized neighborhood prompts, enabling the use of latent node attribute predictability as an anomaly score across various datasets.
Appendix C Additional Experimental Results
C.1. Train the Model on Amazon and Evaluate on Other Datasets
C.2. Sensitivity Analysis
Impact of $\beta$ in anomaly scoring. The performance under different $\beta$ is shown in Fig. 7. As shown in Table 7, the set of $\beta$ values are determined based on the prior we gain from the global average similarity. Specifically, the hyperparameter in anomaly scoring $\beta$ is set to zero by default for graphs with high global similarity such as Amazon and Yelp, and four for those with low global similarity such as Ellitipic and T-Finance, see the detailed global similarity information in Table 1. In the few-shot scoring, $\beta$ is set to $0.5$ by default for graphs with high global similarity. The main reason for these settings is that for graphs with high global average similarity, where node features are relatively similar, the deviations in local neighborhood would be small, and the fluctuations in scoring based on the normal prototype may affect the overall anomaly score. Therefore, we prefer to solely use the abnormal class prototype or assign a smaller weight for the normal class prototype in the anomaly scoring. For nodes with less similar features, we use a mix of normal and abnormal prototypes for anomaly detection. In the few-shot setting, the labeled normal nodes help refine the normal prototypes, allowing the anomaly scores to be computed using both normal and abnormal prototypes.
Figure 7. AUROC and AUPRC of AnomalyGFM w.r.t $\beta$.
Table 7. The value of $\beta$ in anomaly scoring. Sim. indicate the global average edge similarity of the graph.
@@ C.2. Sensitivity Analysis
Impact of commonality dimension $d^{\prime}$. The AUPRC results are shown in Fig. 8. We can see that AnomalyGFM still outperforms all the competing methods on AUPRC under the different dimensions, indicating that AnomalyGFM is generally robust to the feature dimensionality.
Figure 8. AUPRC of AnomalyGFM w.r.t commonality dimension $d^{\prime}$
Table 8. The performance of AnomalyGFM under different initialization hyperparameters for Gaussian distribution used in the class prototypes
Metric Initialization Value Dataset
Reddit Amazon Disney Amazon-all YelpChi-all Tolokers Question Elliptic T-Finance
AUROC $\mu$ = 0, $\sigma$ = 1 0.5974 0.8417 0.6525 0.9032 0.5800 0.5843 0.5280 0.6195 0.5614
$\mu$ = 0, $\sigma$ = 0.5 0.5868 0.8326 0.6172 0.9229 0.5668 0.5981 0.5378 0.6068 0.5902
$\mu$ = 0.1, $\sigma$ = 0.5 0.5779 0.8315 0.7641 0.8756 0.5915 0.6004 0.5437 0.5907 0.5727
$\mu$ = 0.2, $\sigma$ = 1 0.5875 0.8546 0.7218 0.8963 0.5915 0.5975 0.5429 0.6099 0.5922
AUPRC $\mu$ = 0, $\sigma$ = 1 0.0387 0.5790 0.1232 0.6820 0.1813 0.2749 0.0335 0.1333 0.0593
$\mu$ = 0, $\sigma$ = 0.5 0.0378 0.5569 0.1003 0.7334 0.1715 0.2824 0.0345 0.1234 0.0612
$\mu$ = 0.1, $\sigma$ = 0.5 0.0373 0.5485 0.1602 0.6568 0.1890 0.2870 0.0342 0.1096 0.0596
$\mu$ = 0.2, $\sigma$ = 1 0.0380 0.5674 0.1646 0.6959 0.1874 0.2853 0.0339 0.1177 0.0604
Table 9. Training time and inference time (seconds) for different methods.
Methods AnomalyDAE TAM GAT BWGNN UNPrompt AnomalyGFM
Training Time 86.04 479.70 2.43 4.86 2.08 7.45
Inference Time 170.46 252.52 91.88 98.33 19.47 11.00
@@
3:  Initilization of the Gaussian distribution of normal class prototype and abnormal class prototype.
@@
Impact of $\mu$ and $\sigma$ in Gaussian distribution initialization. To evaluate the effectiveness of AnomalyGFM, we set different values for $\mu$ and $\sigma$ during the Gaussian distribution initialization for both the normal class prototype and the abnormal class prototype. The AUROC and AUPRC results are shown in the Table 8. The performance remains stable when we adjust $\mu$ and $\sigma$, indicating that AnomalyGFM is not dependent on the initial Gaussian distribution. This is mainly because the supervised loss function provides sufficient supervision, reducing the reliance on the Gaussian distribution.
Appendix D Time Complexity Analysis
In this section, we analyze the time complexity of pre-training and Fine-tuning of AnomalyGFM. We build the GCN as the backbone of AnomalyGFM to obtain the representation of each node, which takes $O(mdh)$, where $m$ is the number of non-zero elements in matrix $\bf A$ (the number of edge in graph), $d$ is the dimension of representation, and $h$ is the number of feature maps. The computation of the residual feature takes $O(N^{2}d)$. The MLP layer mapping the representation to the anomaly score takes $O(Nd)$. The feature mapping of normal class prototype and abnormal class prototype takes $O(2d^{2})$. The alignment loss between residual features and prototypes take $O(2Nd)$. Thus, the overall complexity of the pre-training of AnomalyGFM is $O(mdh+N^{2}d+3Nd+2d^{2})$. In the inference, we employ the residual feature’s similarity with two prototypes which takes $O(2Nd)$. Therefore, the overall complexity of inference of AnomalyGFM is $O(mdh+N^{2}d+2d^{2}+2Nd)$.
@@ Appendix D Time Complexity Analysis
In Table 9, we report the training time on Facebook and inference time on YelpChi-all of AnomalyGFM. The competing methods, including two representative unsupervised methods AnomalyDAE and TAM, and two supervised methods GAT and BWGNN are used for comparison. Besides, UNPrompt is also included for comparison. The results in Table 9 show that AnomalyGFM requires much less training and inference time compared to TAM and AnomalyDAE, demonstrating the efficiency of the proposed AnomalyGFM. For inference, AnomalyGFM is the most efficient method because anomaly scoring relies solely on measuring the similarity between node representation residuals and prototypes.
Appendix E Algorithms

### v2

@@ 6. Conclusion
Acknowledgments
This research is supported by A*STAR under its MTC YIRG Grant (No. M24N8c0103), the Ministry of Education, Singapore under its Tier-1 Academic Research Fund (No. 24-SIS-SMU-008), the Lee Kong Chian Fellowship, and ARC under Grant DP240101349.
@@ Appendix A Detailed Dataset Description
  • Amazon (Dou et al., 2020): It includes product reviews under the Musical Instrument category. The users with more than 80% of helpful votes were labeled as begin entities, with the users with less than 20% of helpful votes treated as fraudulent entities.
@@
  • Amazon-all (Dou et al., 2020): It includes three types of relations: U-P-U (users reviewing at least one same product), U-S-U (users giving at least one same star rating within one week), and U-V-U (users with top-5% mutual review similarities). Amazon-all is formed by treating the different relations as a single relation following (Chen et al., 2022; Qiao and Pang, 2023).
  • YelpChi-all (Dou et al., 2020): Similar to Amazon-all, YelpChi-all includes three types of edges: R-U-R (reviews posted by the same user), R-S-R (reviews for the same product with the same star rating), and R-T-R (reviews for the same product posted in the same month). YelpChi-all is formed by treating the different relations as a single relation following (Chen et al., 2022; Qiao and Pang, 2023).
@@
  • Question (Platonov et al., 2023): It is collected from the website Yandex Q where the node represents the user and the edge connecting the node represents the user who has answered other’s questions during a one year period. The attribute is the mean of embeddings for words in the description of the user.
  • Elliptic (Weber et al., 2019): It is a Bitcoin transaction network in which the node represents the transactions and the edge is the flow of Bitcoin currency. The Bitcoin transaction is mapped to real-world entities associated with licit categories.
  • T-Finance (Tang et al., 2022): It is a financial transaction network where the node represents an account and the edge represents two accounts that have transaction. The features of each account are related to some attributes of logging like registration days, logging activities, and interaction frequency, etc. The users are labeled as anomalies when they fall into the categories like fraud money laundering and online gambling.
@@
Appendix B Additional Experimental Results
B.1. Few-shot Tuning on Abnormal Nodes
Additionally, tuning the abnormal prototype ${\bf{p}_{a}}$ is also feasible when few-shot abnormal node labels are available. Give some labeled abnormal nodes ${{\mathcal{V}}_{l}^{Test}}$, the fine tuning process is formulated as
(12) ${L_{{{pt}}}}=\sum\limits_{i=1}^{\left|{{\mathcal{V}}_{l}^{Test}}\right|}\left\|{{{\bf{r}}_{i}}-{\bf{p}^{\prime}_{a}}}\right\|_{2}^{2},$
where ${{\bf{p}^{\prime}_{a}}}={{\bf{p}}_{a}}+{g}({{\bf{p}}_{a};\phi})+\Psi_{\mathbf{p}_{a}}$, with a similar formulation applied during the tuning of few-shot normal nodes. To evaluate the performance of AnomalyGFM under few-shot abnormal nodes setting. we include a few labeled abnormal nodes, and treat the unlabeled nodes as normal. These can be combined to serve as the training source. We compare AnomalyGFM with the few-shot GAD method MetaGAD (Xu et al., 2024) and fully supervised methods BWGNN (Tang et al., 2022). We show results in Table 7 for the scenario where few-shot labeled abnormal nodes are available. As shown in Table 7, AnomalyGFM outperforms MetaGAD and BWGNN since they are limited by the scarcity of anomalies on the target graphs, whereas AnomalyGFM has better generalization since it only needs to adapt the pre-trained anomaly prototype with the few anomaly nodes to the target graphs.
Table 7. Comparison under 1/5-shot anomalies
Metric Setting Method Dataset
Amazon-all YelpChi-all
AUROC 1-shot BWGNN (ICML’22) 0.7441 0.5447
MetaGAD (DSAA’24) 0.8602 0.5434
AnomalyGFM 0.9155 0.5941
5-shot BWGNN (ICML’22) 0.8292 0.4759
MetaGAD (DSAA’24) 0.8558 0.5741
AnomalyGFM 0.9246 0.6004
AUPRC 1-shot BWGNN (ICML’22) 0.2947 0.1703
MetaGAD (DSAA’24) 0.7562 0.1681
AnomalyGFM 0.6956 0.1922
5-shot BWGNN (ICML’22) 0.4433 0.1408
MetaGAD (DSAA’24) 0.7567 0.1814
AnomalyGFM 0.7538 0.1935
B.2. Train the Model on Amazon and Evaluate on Other Datasets
Table 8. Comparison of AUROC and AUPRC results on 8 real-world GAD datasets under zero-shot setting with the models trained on Amazon only.
@@ B.2. Train the Model on Amazon and Evaluate on Other Datasets
B.3. Sensitivity Analysis
Impact of $\beta$ in Anomaly Scoring. As shown in Table 9, the set of $\beta$ values are determined based on the prior we gain from the global average similarity in Table 1. It is set to 0 by default for graphs with high global similarity and 4 for graphs with low global similarity. In the few-shot scoring, $\beta$ is set to $0.5$ by default for graphs with high global similarity and four for graphs with low global similarity, allowing the anomaly scores to be computed using both normal and abnormal class prototypes, where the normal class prototype is refined using the labeled normal nodes.
Table 9. The value of $\beta$ in anomaly scoring in both zero-shot and few-shot setting. Sim. indicate the global average edge similarity of the graph.
@@ B.3. Sensitivity Analysis
Impact of $\mu$ and $\sigma$ in Gaussian distribution initialization. To evaluate the effectiveness of AnomalyGFM, we set different values for $\mu$ and $\sigma$ during the Gaussian distribution initialization for both the normal class prototype and the abnormal class prototype, see in the Table 10. The performance remains stable when we adjust $\mu$ and $\sigma$, indicating that AnomalyGFM is not dependent on the initial Gaussian distribution. This is mainly because the supervised loss function provides sufficient supervision, reducing the reliance on the Gaussian distribution.
Table 10. Performance of AnomalyGFM under different initialization hyperparameters for Gaussian distribution
Metric Initialization Dataset
Amazon-all Tolokers Question Elliptic T-Finance
AUROC $\mu$ = 0, $\sigma$ = 1 0.9032 0.5843 0.5280 0.6195 0.5614
$\mu$ = 0, $\sigma$ = 0.5 0.9229 0.5981 0.5378 0.6068 0.5902
$\mu$ = 0.1, $\sigma$ = 0.5 0.8756 0.6004 0.5437 0.5907 0.5727
$\mu$ = 0.2, $\sigma$ = 1 0.8963 0.5975 0.5429 0.6099 0.5922
AUPRC $\mu$ = 0, $\sigma$ = 1 0.6820 0.2749 0.0335 0.1333 0.0593
$\mu$ = 0, $\sigma$ = 0.5 0.7334 0.2824 0.0345 0.1234 0.0612
$\mu$ = 0.1, $\sigma$ = 0.5 0.6568 0.2870 0.0342 0.1096 0.0596
$\mu$ = 0.2, $\sigma$ = 1 0.6959 0.2853 0.0339 0.1177 0.0604
Appendix C Algorithms